OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! simplify the following equation 4x^3 8x^5 ----- divided ------ 9y^4 3y^2

OpenStudy (magbak):

@Mousam @mathmate @agent0smith @uri @Mertsj

OpenStudy (anonymous):

\[\frac{ 1 }{ 6x^{2}y^{2} }\]

OpenStudy (magbak):

That is wrong

OpenStudy (anonymous):

???

OpenStudy (magbak):

This are the answer choices and C is wrong 6x^2y^2 6 ------ x^2y^2 6x^2y^2 x^2y^2 ------- 6

OpenStudy (anonymous):

\[\frac{ 4x^{3} }{ 9y^{4}} \frac{ 3y^{2} }{ 8x^{5} }\]

OpenStudy (anonymous):

\[\frac{ 12 }{ 72x^{2}y^{2} }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 6x^{2}y^{2} }\]

OpenStudy (anonymous):

i dunno where i've gone wrong... do you see it?

OpenStudy (magbak):

Me to I do not know

OpenStudy (magbak):

@agent0smith

OpenStudy (anonymous):

can you recheck the choices please...

OpenStudy (magbak):

yes sure they are 6x^2y^2 6 ------ x^2y^2 6x^2y^2 x^2y^2 ------- 6

OpenStudy (anonymous):

wow... i acutally have no idea what is wrong....

OpenStudy (magbak):

@zscdragon

OpenStudy (agent0smith):

Is the first not 1/6x^2y^2? @Mousam appears correct.

OpenStudy (magbak):

@Mertsj @zepdrix @Luigi0210

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yea.. because the first and third choices are the same....

OpenStudy (magbak):

Sorry is is 6x^2y^2 6 ------ x^2y^2 1 ------- 6x^2y^2 x^2y^2 ------- 6

OpenStudy (anonymous):

haha... it is the third choice

OpenStudy (anonymous):

i was worried for a moment...

OpenStudy (magbak):

No but the third one is wrong because I enter that as my choice and I got it wrong

OpenStudy (magbak):

@agent0smith

OpenStudy (anonymous):

???

OpenStudy (magbak):

I do not know

OpenStudy (magbak):

@JA1

OpenStudy (magbak):

@AccessDenied

OpenStudy (magbak):

@kropot72 @robtobey @terenzreignz @radar @timo86m @zscdragon

OpenStudy (agent0smith):

The third one is correct. Tagging more people isn't going to make it any more correct.

OpenStudy (magbak):

Yes but the teacher marked it worong like 5 times and told me it is wrong

OpenStudy (agent0smith):

Then your teacher is wrong or the problem is wrong 4x^3 8x^5 ----- divided ------ 9y^4 3y^2\[\Large \frac{ 4x^3* 3y^2 }{ 9y^4 *8x^5 } = \]\[\Large \frac{ 12 x^3 y^2 }{ 72 x^5 y^4 } = \frac{ 1 }{6x^2 y^2 }\]

OpenStudy (anonymous):

you should go and slap your teacher... lol

OpenStudy (magbak):

Not proof it was just marked wrong and I believe you. @Loser66

OpenStudy (magbak):

My teacher is texting me the answer please wait for verification

OpenStudy (agent0smith):

No let's not wait for verification, the answer has been verified by wolfram freakin alpha!

OpenStudy (magbak):

Hey NOT MY FAULT SO DO NOT BE RUDE OK

OpenStudy (agent0smith):

lol it was a joke...

OpenStudy (loser66):

@magbak, I think your teacher tests whether you are sure about your answer or not. Be confident, please

OpenStudy (magbak):

ok

OpenStudy (magbak):

Ok I have 3 or 2 more question please help me I am sorry. I am not being hardheaded.